Watson, Inc., is an all-equity firm. The cost of the company’s equity is currently 11 percent, and the risk-free rate is 5.1 percent. The company is currently considering a project that will cost $11.88 million and last six years. The company uses straight-line depreciation. The project will generate revenues minus expenses each year in the amount of $3.36 million.

  

If the company has a tax rate of 35 percent, what is the net present value of the project? (Enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  Net present value $
Answer #1

Cash flows from Year 1 to Year 6 = Incremental earnings * (1 - tax rate) + Depreciation * Tax rate

Incremental earnings = 3.36 million

Depreciation = 11.88 / 6 = 1.98 million

Cash flow = 3.36 * (1 - 0.35) + 1.98 * 0.35

= 2.877 Million

PV of cash flows = Cash flow * PVIFA (n,i) where n =6 and i = 11%

= 2.877 * (1 - (1+11%)^-6) / 0.11

= 12.17126 million

NPV = PV of cash flows - Initial investment = 12.17126 - 11.88

= $ 291260